# FedDiv: Collaborative Noise Filtering for Federated Learning with Noisy Labels

### 19 Dec 2023 v1

@@ Introduction
Figure 2: An overview of the training procedure proposed in FedDiv. In this work, the parameters of a local neural model and a local noise filter are simultaneously learned on each client during the local training sessions, while both types of parameters are aggregated on the server.
@@
  • We propose FedDiv, a novel one-stage framework for federated learning with noisy labels. To enable stable training, FedDiv learns a global noise filter by distilling the complementary knowledge from all clients while performing label noise filtering locally on every client.
@@
Method Best Test Accuracy $\pm$ Standard Deviation
$\rho$=0.4 $\rho$=0.6 $\rho$=0.8 $\rho$=1.0
$\tau$=0.0 $\tau$=0.5 $\tau$=0.0 $\tau$=0.5 $\tau$=0.0 $\tau$=0.5 $\tau$=0.0 $\tau$=0.5
FedAvg (McMahan et al. 2017) 89.46$\pm$0.39 88.31$\pm$0.80 86.09$\pm$0.50 81.22$\pm$1.72 82.91$\pm$1.35 72.00$\pm$2.76 - -
FedProx (Li et al. 2020) 88.54$\pm$0.33 88.20$\pm$0.63 85.80$\pm$0.41 85.25$\pm$1.02 84.17$\pm$0.77 80.59$\pm$1.49 - -
RoFL (Yang et al. 2022b) 88.25$\pm$0.33 87.20$\pm$0.26 87.77$\pm$0.83 83.40$\pm$1.20 87.08$\pm$0.65 74.13$\pm$3.90 - -
ARFL (Fu et al. 2019) 85.87$\pm$1.85 83.14$\pm$3.45 76.77$\pm$1.90 64.31$\pm$3.73 73.22$\pm$1.48 53.23$\pm$1.67 - -
JointOpt (Tanaka et al. 2018a) 84.42$\pm$0.70 83.01$\pm$0.88 80.82$\pm$1.19 74.09$\pm$1.43 76.13$\pm$1.15 66.16$\pm$1.71 - -
DivideMix (Li, Socher, and Hoi 2019) 77.35$\pm$0.20 74.40$\pm$2.69 72.67$\pm$3.39 72.83$\pm$0.30 68.66$\pm$0.51 68.04$\pm$1.38 - -
[…]
FedDiv (Ours) 94.42$\pm$0.29 94.30$\pm$0.19 93.67$\pm$0.22 93.41$\pm$0.21 92.98$\pm$0.60 91.44$\pm$0.25 91.20$\pm$0.40 86.51$\pm$3.18
[…]
@@
FedAvg (McMahan et al. 2017) 78.88$\pm$2.34 75.98$\pm$2.92 67.75$\pm$4.38
FedProx (Li et al. 2020) 83.32$\pm$0.98 80.40$\pm$0.94 73.86$\pm$2.41
RoFL (Yang et al. 2022b) 79.56$\pm$1.39 72.75$\pm$2.21 60.72$\pm$3.23
ARFL (Fu et al. 2019) 60.19$\pm$3.33 55.86$\pm$3.30 45.78$\pm$2.84
JointOpt (Tanaka et al. 2018a) 72.19$\pm$1.59 66.92$\pm$1.89 58.08$\pm$2.18
DivideMix (Li, Socher, and Hoi 2019) 65.70$\pm$0.35 61.68$\pm$0.56 56.67$\pm$1.73
FedCorr (Xu et al. 2022) 90.52$\pm$0.89 88.03$\pm$1.08 81.57$\pm$3.68
@@
$\begin{split}\textbf{\text{p}}(&\text{``clean''}|x,y;{\theta}_{k}^{(t)})=P(z=1|x,y;{\theta}_{k}^{(t)}).\end{split}$ (6)
@@
Method Best Test Accuracy $\pm$ Standard Deviation
$\rho$=0.4 $\rho$=0.6 $\rho$=0.8 $\rho$=1.0
$\tau$=0.5 $\tau$=0.5 $\tau$=0.5 $\tau$=0.5
FedAvg (McMahan et al. 2017) 64.41$\pm$1.79 53.51$\pm$2.85 44.45$\pm$2.86 -
FedProx (Li et al. 2020) 65.09$\pm$1.46 57.51$\pm$2.01 51.24$\pm$1.60 -
RoFL (Yang et al. 2022b) 59.42$\pm$2.69 46.24$\pm$3.59 36.65$\pm$3.36 -
ARFL (Fu et al. 2019) 51.53$\pm$4.38 33.03$\pm$1.81 27.47$\pm$1.08 -
JointOpt (Tanaka et al. 2018a) 58.43$\pm$1.88 44.54$\pm$2.87 35.25$\pm$3.02 -
DivideMix (Li, Socher, and Hoi 2019) 43.25$\pm$1.01 40.72$\pm$1.41 38.91$\pm$1.25 -
[…]
FedDiv (Ours) 74.86$\pm$0.91 72.37$\pm$1.12 65.49$\pm$2.20 45.95$\pm$5.36
Method\$(p,\alpha_{Dir})$ $(0.7,10)$
FedAvg (McMahan et al. 2017) 64.75$\pm$1.75
FedProx (Li et al. 2020) 65.72$\pm$1.30
RoFL (Yang et al. 2022b) 59.31$\pm$4.14
ARFL (Fu et al. 2019) 48.03$\pm$4.39
JointOpt (Tanaka et al. 2018a) 59.84$\pm$1.99
DivideMix (Li, Socher, and Hoi 2019) 39.76$\pm$1.18
FedCorr (Xu et al. 2022) 72.73$\pm$1.02
FedDiv (Ours) 74.47$\pm$0.34
Table 4: Best test accuracy (%) of FedDiv and existing SOTA methods on CIFAR-100 with non-IID setting at the noise level $(\rho,\tau)=(4.0,0)$.
RoFL (Yang et al. 2022b) ARFL (Fu et al. 2019) JointOpt (Tanaka et al. 2018a) Dividemix (Li, Socher, and Hoi 2019) FedCorr (Xu et al. 2022) FedDiv (Ours)
70.39 70.91 71.78 68.83 72.55 72.96$\pm$0.43
Table 5: Best test accuracy (%) of FedDiv and existing SOTA methods on Clothing1M with non-IID setting.
@@
$\begin{split}~{}\bar{\mathcal{D}}_{k}\leftarrow\{\left(x,y\right)|\hat{y}(x)=\tilde{y}(x),\forall(x,y)\in\mathcal{D}_{k}^{\text{clean}}\cup\mathcal{D}_{k}^{\text{relab}}\}.\end{split}$ (10)
We update the training dataset for further optimizing the local model as follows,
$\hat{\mathcal{D}}_{k}=\begin{cases}\bar{\mathcal{D}}_{k},&\text{if}~{}\hat{\delta}_{k}\geq 0.1,\\ \mathcal{D}_{k},&\text{if}~{}\hat{\delta}_{k}<0.1.\end{cases}$ (11)
Once having the optimized local model $\theta^{(t)}_{k}$ from a local training session, we use it to update $\hat{p}_{k}$ with momentum as follows,
@@
To be fair, we here adopt the consistent experimental setups with FedCorr (Xu et al. 2022) to assess the efficacy of our proposed approach FedDiv. Further details, e.g., data partitions, implementation, and baseline models, are provided in the supplementary document (abbr. Supp; see below), which can also be found at https://github.com/lijichang/FLNL-FedDiv/blob/main/supp.pdf.
Datasets and data partitions. We validate FedDiv’s superiority on three classic benchmark datasets, including two synthetic datasets namely CIFAR-10 (Krizhevsky 2009) and CIFAR-100 (Krizhevsky 2009), and one real-world noisy dataset, i.e., Clothing1M (Xiao et al. 2015). Like (Xu et al. 2022), we take both IID and non-IID data partitions into account on CIFAR-10 and CIFAR-100, but only consider non-IID data partitions on Clothing1M. Under the IID data partitions, each client is randomly assigned the same number of samples with respect to each class. For non-IID data partitions, it is constructed using a Dirichlet distribution (Lin et al. 2020) with two pre-defined parameters, namely the fixed probability $p$ and the concentration parameter $\alpha_{Dir}$; See Supp for more details.
Label noise settings. Similar to (Xu et al. 2022), all clients are separated into a subset of clean clients with only clean samples and a subset of noisy clients potentially containing samples with label noise at varying degrees. The noise level, ${\delta}_{k}$, for the $k$-th client is defined as follows:
@@
Implementation. We set $\omega$ and $\mathcal{T}$ to 950, 900 and 100, and 0.1, 0.1 and 0.02 on CIFAR-10, CIFAR-100 and Clothing1M, respectively, while we also set the confidence threshold $\zeta=0.75$ for relabeling on all datasets. Note that, to enable faster convergence, we warm up local neural network models of each client for $\mathcal{T}_{wu}$ iterations (not training rounds; see (Xu et al. 2022)) using the standard cross-entropy loss. Additionally, to be fair, most of our implementation details involving both local training and model aggregation are consistent with FedCorr (Xu et al. 2022) for each dataset under all federated settings and all label noise settings. For more implementation details, please see Supp.
@@
Figure 3: The accuracy of noisy label identification across different clients. The lightblue dotted line represents the actual noise level of each client, while the brightly colored dotted lines indicate the noise filtering performance w.r.t different noise filters. The experiment is conducted on CIFAR-100 with the non-IID data partition under the noise setting $(\rho,\tau)=(0.4,0.0)$. (Best viewed zoomed in.)
Comparison with State-of-the-Arts
Tables 1-5 summarize classification performance of FedDiv against state-of-the-art (SOTA) F-LNL methods on CIFAR-10, CIFAR-100, and Clothing1M across various noise settings for both IID and non-IID data partitions. Comparison results, based on mean accuracy and standard deviation over five trials, demonstrate FedDiv’s significant superiority over existing F-LNL algorithms, especially in challenging cases. For instance, in IID data partitions, Table 3 illustrates FedDiv outperforming FedCorr on CIFAR-100 by 6.39% and 16.46% in the toughest noise settings: $(\rho,\tau)=(0.8,0.5)$ and $(\rho,\tau)=(1.0,0.5)$. Similarly, for non-IID partitions in Table 2, FedDiv consistently surpasses FedCorr by 3.74% in the most challenging setting $(p,\alpha_{Dir})=(0.3,10)$ on CIFAR-10. Additionally, in Table 5, FedDiv exhibits a 0.41% improvement over FedCorr on Clothing1M, indicating its efficacy in real-world label noise distributions.
@@
Noise level ($\rho$, $\tau$)
(0.8, 0.0)
(0.6, 0.5)
(0.8, 0.5)
(0.4,0.0)
Method\($p$, $\alpha_{Dir}$)
$\dagger$
(0.3, 10)
$\dagger$
(0.7, 10)
FedCorr (Xu et al. 2022)
91.52$\pm$0.50
81.57$\pm$3.68
59.10$\pm$5.12
72.73$\pm$1.02
FedDiv (Ours)
92.98$\pm$0.60
85.31$\pm$2.28
65.49$\pm$2.20
74.47$\pm$0.34
FedDiv (Degraded)
92.31$\pm$0.61
83.22$\pm$2.61
62.23$\pm$2.37
73.06$\pm$0.93
FedDiv (Local filter)
91.14$\pm$0.48
81.34$\pm$3.65
60.97$\pm$1.65
71.37$\pm$0.76
FedDiv w/o Relab. & w/o PCS
91.43$\pm$0.87
82.17$\pm$3.06
61.45$\pm$3.43
73.09$\pm$1.89
FedDiv w/o PCS
92.29$\pm$0.96
82.83$\pm$2.59
62.67$\pm$2.51
73.66$\pm$0.96
FedDiv w/o PCS.debias
92.54$\pm$0.72
84.19$\pm$3.31
64.18$\pm$2.89
74.02$\pm$0.65
FedDiv w/o $\mathcal{L}_{reg}$
83.60$\pm$3.65
72.43$\pm$1.29
Table 6: Ablation study results for CIFAR-10 and CIFAR-100, in varying noise levels and both IID and non-IID data partition settings, with $\dagger$ indicating the IID data partitions.
@@
To underscore the efficacy of FedDiv, we perform an ablation study to demonstrate the effect of each component. More analysis results can be found in Supp.
@@ Conclusions
In this paper, we have presented FedDiv to deal with federated learning with noisy labels (F-LNL). It can effectively respond to the challenges in F-LNL tasks involving both data heterogeneity and noise heterogeneity while taking privacy concerns into account. On the basis of an FL framework, we first propose Federated Noise Filtering to separate clean samples from noisy ones on each client, thereby diminishing the instability during training. Then we perform relabeling to assign pseudo-labels to noisy samples with high predicted confidence. In addition, we introduce a Predictive Consistency based Sampler to identify credible local data for local model training, thus avoiding label noise memorization and further improving training stability. Experiments as well as comprehensive ablation analysis have revealed FedDiv’s superiority in handling F-LNL tasks.
@@ Acknowledgements
Collection of randomly selected clients at current round
@@
Current index for the client randomly selected from $\mathcal{M}$ or $\mathcal{S}$
@@
More reliable labeled samples re-selected by PCS
@@
Actual noise level at client $k$
@@
Estimated noise level at client $k$
@@
Global network parameters at round $t$
@@
Local network parameters being optimized during local training for client $k$ at round $t$
@@
Local network parameters for client $k$ at round $t$
@@
16:           Obtain $\hat{\mathcal{D}}_{k}$ using Eqs. (9), (10), (11)
@@
Baselines. We compare FedDiv with existing state-of-the-art (SOTA) F-LNL methods, including FedAvg (McMahan et al. 2017), FedProx (Li et al. 2020), RoFL (Yang et al. 2022b), ARFL (Yang et al. 2022b), JointOpt (Tanaka et al. 2018a), DivideMix (Li, Socher, and Hoi 2019), FedCorr (Xu et al. 2022), borrowing the reported experimental results from (Xu et al. 2022). Specifically, FedAvg and FedProx are both classic FL algorithms, while RoFL, ARFL and FedCorr are three existing F-LNL methods. Besides, DivideMix and JointOpt are two exiting representative C-LNL algorithms, which are adapted by FedCorr into the FL scenarios.
How to set $\omega$, $\mathcal{T}$ and $\mathcal{T}_{wu}$. In this work, we streamlined the multi-stage F-LNL process proposed in FedCorr (Xu et al. 2022) into a one-stage process, avoiding the complexity of executing multiple intricate steps across different stages as in FedCorr. However, for fair comparisons, we maintained an equivalent number of communication rounds as in FedCorr. This totals $\mathcal{T}_{wu}$, encompassing federated pre-processing from FedCorr’s training iterations, and $\mathcal{T}$, covering both federated fine-tuning and usual training stages involving FedCorr. Notably, within $\mathcal{T}_{wu}$, we solely utilize standard cross-entropy loss to warm up the local neural network models for faster convergence.

### 16 Feb 2024 v3

@@ Introduction
Figure 2: An overview of the training procedure proposed by FedDiv. In this work, the parameters of a local neural model and a local noise filter are simultaneously learned on each client during the local training sessions, while both types of parameters are aggregated on the server.
@@
  • We propose a novel one-stage framework, FedDiv, for addressing the task of Federated Learning with Noisy Labels (F-LNL). To enable stable training, FedDiv learns a global noise filter by distilling the complementary knowledge from all clients while performing label noise filtering locally on every client.
@@
Method Best Test Accuracy $\pm$ Standard Deviation
$\rho$=0.4 $\rho$=0.6 $\rho$=0.8
$\tau$=0.0 $\tau$=0.5 $\tau$=0.0 $\tau$=0.5 $\tau$=0.0 $\tau$=0.5
FedAvg 89.46$\pm$0.39 88.31$\pm$0.80 86.09$\pm$0.50 81.22$\pm$1.72 82.91$\pm$1.35 72.00$\pm$2.76
RoFL 88.25$\pm$0.33 87.20$\pm$0.26 87.77$\pm$0.83 83.40$\pm$1.20 87.08$\pm$0.65 74.13$\pm$3.90
ARFL 85.87$\pm$1.85 83.14$\pm$3.45 76.77$\pm$1.90 64.31$\pm$3.73 73.22$\pm$1.48 53.23$\pm$1.67
JointOpt 84.42$\pm$0.70 83.01$\pm$0.88 80.82$\pm$1.19 74.09$\pm$1.43 76.13$\pm$1.15 66.16$\pm$1.71
DivideMix 77.35$\pm$0.20 74.40$\pm$2.69 72.67$\pm$3.39 72.83$\pm$0.30 68.66$\pm$0.51 68.04$\pm$1.38
FedCorr 94.01$\pm$0.22 94.15$\pm$0.18 92.93$\pm$0.25 92.50$\pm$0.28 91.52$\pm$0.50 90.59$\pm$0.70
FedDiv (Ours) 94.42$\pm$0.29 94.30$\pm$0.19 93.67$\pm$0.22 93.41$\pm$0.21 92.98$\pm$0.60 91.44$\pm$0.25
Table 1: Best test accuracy (%) of FedDiv and existing SOTA methods on CIFAR-10 with IID setting at diverse noise levels.
@@
FedAvg 78.88$\pm$2.34 75.98$\pm$2.92 67.75$\pm$4.38
RoFL 79.56$\pm$1.39 72.75$\pm$2.21 60.72$\pm$3.23
ARFL 60.19$\pm$3.33 55.86$\pm$3.30 45.78$\pm$2.84
JointOpt 72.19$\pm$1.59 66.92$\pm$1.89 58.08$\pm$2.18
DivideMix 65.70$\pm$0.35 61.68$\pm$0.56 56.67$\pm$1.73
FedCorr 90.52$\pm$0.89 88.03$\pm$1.08 81.57$\pm$3.68
@@
$\begin{split}\textbf{\text{p}}(&\text{``clean''}|x,y;{\theta}^{(t)})=P(z=1|x,y;{\theta}^{(t)}).\end{split}$ (6)
@@
Method Best Test Accuracy $\pm$ Standard Deviation
$\rho$=0.4 $\rho$=0.6 $\rho$=0.8
$\tau$=0.5 $\tau$=0.5 $\tau$=0.5
FedAvg 64.41$\pm$1.79 53.51$\pm$2.85 44.45$\pm$2.86
RoFL 59.42$\pm$2.69 46.24$\pm$3.59 36.65$\pm$3.36
ARFL 51.53$\pm$4.38 33.03$\pm$1.81 27.47$\pm$1.08
JointOpt 58.43$\pm$1.88 44.54$\pm$2.87 35.25$\pm$3.02
DivideMix 43.25$\pm$1.01 40.72$\pm$1.41 38.91$\pm$1.25
FedCorr 74.43$\pm$0.72 66.78$\pm$4.65 59.10$\pm$5.12
FedDiv (Ours) 74.86$\pm$0.91 72.37$\pm$1.12 65.49$\pm$2.20
@@
Dataset CIFAR-100 Clothing1M
Noise level ($\rho$, $\tau$) (0.4, 0.0) -
Method\($p$, $\alpha_{Dir}$) (0.7, 10) -
FedAvg 64.75$\pm$1.75 70.49
RoFL 59.31$\pm$4.14 70.39
ARFL 48.03$\pm$4.39 70.91
JointOpt 59.84$\pm$1.99 71.78
DivideMix 39.76$\pm$1.18 68.83
FedCorr 72.73$\pm$1.02 72.55
FedDiv (Ours) 74.47$\pm$0.34 72.96$\pm$0.43
Table 4: Best test accuracy (%) of FedDiv and existing SOTA methods on CIFAR-100 and Clothing1M under the non-IID data partitions.
@@
$\begin{split}~{}\mathcal{D}_{k}^{\text{resel}}\leftarrow\{\left(x,y\right)|\hat{y}(x)=\tilde{y}(x),\forall(x,y)\in\mathcal{D}_{k}^{\text{clean}}\cup\mathcal{D}_{k}^{\text{relab}}\}.\end{split}$ (10)
Similar to (Xu et al. 2022), we update the training dataset for further optimizing the local model as follows,
$\hat{\mathcal{D}}_{k}=\begin{cases}\mathcal{D}_{k}^{\text{resel}},&\text{if}~{}\hat{\delta}_{k}\geq 0.1,\\ \mathcal{D}_{k},&\text{if}~{}\hat{\delta}_{k}<0.1.\end{cases}$ (11)
Once having the optimized $\theta^{(t)}_{k}$ from a local training session, we use it to update $\hat{p}_{k}$ with momentum as follows,
@@
To be fair, we here adopt the consistent experimental setups with FedCorr (Xu et al. 2022) to assess the efficacy of our proposed approach FedDiv. Further details, e.g., data partitions and additional implementations and analysis, are provided in the supplementary document (abbr. Supp [^1] https://github.com/lijichang/FLNL-FedDiv/blob/main/supp.pdf).
Datasets and data partitions. We validate FedDiv’s superiority on three classic benchmark datasets, including two synthetic datasets namely CIFAR-10 (Krizhevsky 2009) and CIFAR-100 (Krizhevsky 2009), and one real-world noisy dataset, i.e., Clothing1M (Xiao et al. 2015). Like (Xu et al. 2022), we take both IID and non-IID data partitions into account on CIFAR-10 and CIFAR-100, but only consider non-IID data partitions on Clothing1M. Under the IID data partitions, each client is randomly assigned the same number of samples with respect to each class. For non-IID data partitions, it is constructed using a Dirichlet distribution (Lin et al. 2020) with two pre-defined parameters, namely the fixed probability $p$ and the concentration parameter $\alpha_{Dir}$.
Label noise settings. Similar to (Xu et al. 2022), the noise level for the $k$-th client can be defined as follows:
@@
Figure 3: The accuracy of noisy label identification across different clients, with the experiment conducted on CIFAR-100 with the non-IID data partition under the noise setting $(\rho,\tau)=(0.4,0.0)$. (Best viewed zoomed in.)
Baselines. We compare FedDiv with existing state-of-the-art (SOTA) F-LNL methods, including FedAvg (McMahan et al. 2017), RoFL (Yang et al. 2022b), ARFL (Yang et al. 2022b), JointOpt (Tanaka et al. 2018a), DivideMix (Li, Socher, and Hoi 2019), and FedCorr (Xu et al. 2022). Their experimental results reported in this paper are borrowed from (Xu et al. 2022).
Implementations. We set $\omega$ and $\mathcal{T}$ to 950, 900 and 100, and 0.1, 0.1 and 0.02 on CIFAR-10, CIFAR-100 and Clothing1M, respectively, while we also set the confidence threshold $\zeta=0.75$ for relabeling on all datasets. Note that, to enable faster convergence, we warm up local neural network models of each client for $\mathcal{T}_{wu}$ iterations (not training rounds; see (Xu et al. 2022)) using MixUp regularization (Zhang et al. 2018). Additionally, to be fair, most of our implementation details involving both local training and model aggregation are consistent with FedCorr (Xu et al. 2022) for each dataset under all federated settings and all label noise settings.
@@
Comparisons with State-of-the-Arts
Tables 1-4 summarize classification performance of FedDiv against state-of-the-art (SOTA) F-LNL methods on CIFAR-10, CIFAR-100, and Clothing1M across various noise settings for both IID and non-IID data partitions. Comparison results, based on mean accuracy and standard deviation over five trials, demonstrate FedDiv’s significant superiority over existing F-LNL algorithms, especially in challenging cases. For instance, in IID data partitions, Table 3 illustrates FedDiv outperforming FedCorr on CIFAR-100 by 6.39% in the toughest noise setting with $(\rho,\tau)=(0.8,0.5)$. Similarly, for non-IID partitions in Table 2, FedDiv consistently surpasses FedCorr by 3.74% in the most challenging setting $(p,\alpha_{Dir})=(0.3,10)$ on CIFAR-10. Additionally, in Table 4, FedDiv exhibits a 0.41% improvement over FedCorr on Clothing1M, indicating its efficacy in real-world label noise distributions.
@@
Noise level ($\rho$, $\tau$) (0.6, 0.5) (0.4,0.0)
Method\($p$, $\alpha_{Dir}$) (0.3, 10) (0.7, 10)
FedDiv (Ours) 85.31$\pm$2.28 74.47$\pm$0.34
FedDiv (Degraded) 83.22$\pm$2.61 73.06$\pm$0.93
FedDiv (Local filter) 81.34$\pm$3.65 71.37$\pm$0.76
FedDiv w/o Relab. & w/o PCS 82.17$\pm$3.06 73.09$\pm$1.89
FedDiv w/o PCS 82.83$\pm$2.59 73.66$\pm$0.96
FedDiv w/o $\mathcal{L}_{reg}$ 83.60$\pm$3.65 72.43$\pm$1.29
Table 5: Ablation study results for CIFAR-10 and CIFAR-100, with varying noise levels under non-IID data partitions.
@@
To underscore the efficacy of FedDiv, we perform an ablation study to demonstrate the effect of each component.
@@ Conclusions
In this paper, we have presented FedDiv to handle the task of Federated Learning with Noisy Labels (F-LNL). It can effectively respond to the challenges in F-LNL tasks involving both data heterogeneity and noise heterogeneity while taking privacy concerns into account. On the basis of an FL framework, we first propose Federated Noise Filtering to separate clean samples from noisy ones on each client, thereby diminishing the instability during training. Then we perform relabeling to assign pseudo-labels to noisy samples with high predicted confidence. In addition, we introduce a Predictive Consistency based Sampler to identify credible local data for local model training, thus avoiding label noise memorization and further improving training stability. Experiments as well as comprehensive ablation analysis have revealed FedDiv’s superiority in handling F-LNL tasks.
@@ Acknowledgements
Collection of clients randomly selected at current round
@@
Current index for the client selected from $\mathcal{M}$ or $\mathcal{S}$
@@
$\mathcal{D}_{k}^{\text{resel}}$
More reliable labeled samples re-selected by PCS
@@
Training data for further optimization of the local model
@@
Actual noise level for client $k$
@@
Estimated noise level for client $k$
@@
Global network model at round $t$
@@
Local network model being optimized during local training for client $k$ at round $t$
@@
Local network model for client $k$ at round $t$
@@
16:           Obtain $\mathcal{D}_{k}^{\text{opt}}$ using Eqs. (9), (10), (11)
@@
How to set $\omega$, $\mathcal{T}$ and $\mathcal{T}_{wu}$. In this work, we streamlined the multi-stage F-LNL process proposed in FedCorr (Xu et al. 2022) into a one-stage process, avoiding the complexity of executing multiple intricate steps across different stages as in FedCorr. However, for fair comparisons, we maintained an equivalent number of communication rounds as in FedCorr. This totals $\mathcal{T}_{wu}$, encompassing federated pre-processing from FedCorr’s training iterations, and $\mathcal{T}$, covering both federated fine-tuning and usual training stages involving FedCorr. Notably, within $\mathcal{T}_{wu}$, we solely utilize MixUp regularization (Zhang et al. 2018) to warm up the local neural network models for faster convergence.